OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (oaktree):

So, first use the two points to find the slope.

OpenStudy (oaktree):

Slope is\[(y _{2} - y_{1})\div(x_{2}-x_{1})\]

OpenStudy (oaktree):

So, 13+11/2+4 = 24/6=4.

OpenStudy (oaktree):

So, the line is y=4x+b, but we need the y-intercept. To do this, we plug in a point.

OpenStudy (oaktree):

Plugging in (-5,3), we have 3=4(-5)+b, or b=23.

OpenStudy (oaktree):

So the line is y=4x+23.

OpenStudy (anonymous):

okay so the slop would be 4 correct?